how would i write (-3,9), m=8 in slope intercept form?
Answer by nerdybill(7384)

how would i write (-3,9), m=8 in slope intercept form?
.
Plug given information into the "point-slope" form of a line:
y - y1 = m(x - x1)
y - 9 = 8(x - (-3))
y - 9 = 8(x + 3)
y - 9 = 8x + 24
y = 8x + 33 (this is what they're looking for)